# X2I: Seamless Integration of Multimodal Understanding into Diffusion Transformer via Attention Distillation

### v2

@@ Abstract
Text-to-image (T2I) models are well known for their ability to produce highly realistic images, while multimodal large language models (MLLMs) are renowned for their proficiency in understanding and integrating multiple modalities. However, currently there is no straightforward and efficient framework to transfer the multimodal comprehension abilities of MLLMs to T2I models to enable them to understand multimodal inputs. In this paper, we propose the X2I framework, which endows Diffusion Transformer (DiT) models with the capability to comprehend various modalities, including multilingual text, screenshot documents, images, videos, and audio. X2I is trained using merely 100K English corpus with 160 GPU hours. Building on the DiT teacher model, we adopt an innovative distillation method to extract the inference capabilities of the teacher model and design a lightweight AlignNet structure to serve as an intermediate bridge. Compared to the teacher model, X2I shows a decrease in performance degradation of less than 1% while gaining various multimodal understanding abilities, including multilingual to image, image to image, image-text to image, video to image, audio to image, and utilizing creative fusion to enhance imagery. Furthermore, it is applicable for LoRA training in the context of image-text to image generation, filling a void in the industry in this area. We further design a simple LightControl to enhance the fidelity of instructional image editing. Finally, extensive experiments demonstrate the effectiveness, efficiency, multifunctionality, and transferability of our X2I. The open-source code and checkpoints for X2I can be found at the following link: https://github.com/OPPO-Mente-Lab/X2I.
footnotetext: 🖂Corresponding authors.
@@ 1 Introduction
The text encoder in T2I is critical for generating semantically relevant images, as seen from early models such as CLIP [59] to later models such as T5 [60] and advanced LLMs [74, 13]. The main goal is to improve the semantic understanding of text without considering the impact of other modalities on visual output. In contrast, VLMs, including early approaches like Flamingo [2] and BLIP2 [35], along with newer models like Llama-Vision [12], QwenVL [3], and InternVL [10], have effectively integrated pre-trained visual encoders with LLMs for enhanced text and visual comprehension. Is there a more streamlined way to transfer VLM or MLLM capabilities to DiT to enable diverse modal inputs? The industry is primarily tackling this challenge through two strategies. One method, similar to PEA-Diffusion [46], utilizes feature distillation by aligning multilingual LLM encoders to the original dimensions using MLP. However, it requires substantial resources of 1600 GPU hours and large-scale image-text datasets. This distillation relies on the U-Net[65] block’s outputs and limits multimodal input support. Another method is the VLMs-based replacement of T2I [44], but it is costly, requiring extensive training or fine-tune of T2I or MLLMs, necessitating substantial data quality and quantity [70, 22]. Alternatively, a solution akin to GlueGen[57] could be employed to align the features of a single-modal or multimodal encoder with the latent space of existing T2I models. However, in terms of strategy, aligning solely through the encoder still results in suboptimal overall performance.
@@
In this work, we introduce an efficient framework, X2I, along with a simple LightControl to facilitate the transition enhancement of images from weak fidelity to strong high fidelity. Utilizing a DiT-based image generation model as a teacher, the student model mirrors the entire architecture of the teacher model but substitutes its text encoder with MLLMs. Both models are capable of processing the same linguistic input. To optimize our training data without gathering images, we opt to distill the teacher model’s inference abilities. In T2I models, this necessitates performing distillation during the reverse denoising phase. Furthermore, we rapidly and effectively impart the teacher model’s visual generation skills to the student model by ingeniously crafting the AlignNet structure and distilling the attention. As a result, post-alignment, the student model acquires the multimodal comprehension abilities of MLLMs during image generation. Additionally, we formulate a template input format for the student model to efficiently interpret instructions across various modalities. Recognizing that MLLMs typically excel at grasping the global semantic features of visual content, we create LightControl to extract structured data from reference images, thereby boosting visual fidelity and allowing for precise and controllable image editing.
@@
After training the X2I, we conduct various experiments to compare the capabilities of T2I and instrutional-based image editing. Additionally, we conduct subjective experiments for image generation with various multimodal mixed inputs, including multilingual, image (I2I), image-text (IT2I), video (V2I), and audio (A2I). The results demonstrate that X2I acquires robust multimodal understanding capabilities with less than 1% performance degradation in the T2I generation. Furthermore, for more personalized tasks like I2I or IT2I, we conduct experiments involving the training of traditional LoRA [23] to fulfill industry requirements, addressing a training gap in LoRA for these purposes. Moreover, to enhance the adaptability of the framework, only the AlignNet parameters within the model are updated, enabling the aligned student model to accommodate a variety of downstream tasks, such as LoRA, ControlNet [72], IP-Adapter [88], different fine-tuning models, and compression models. Notably, some objective metrics for tasks involving ControlNet and IP-Adapter show improvements over the teacher model. This improvement is due to X2I’s robust support for image conditional information input, which facilitates further extraction of image features. We also perform experimental analyses on training efficiency, attaining 98.2% of the teacher model’s performance with 48K training samples.
In summary, our key contributions are:
@@ 2 Related Work
Diffusion Models and Text Encoder. Diffusion models have shown immense potential in the field of T2I. Earlier work [67, 64, 17, 86] uses U-Net as the backbone of the diffusion model, which, to some extent, impacted the scalability of the models. Inspired by DiT [54], many works [8, 32, 17, 36] replace the U-Net backbone with transformers, greatly improving the quality of the generated images. In order to improve the understanding of the diffusion model’s prompts, Stable Diffusion (SD) used CLIP to encode text information, while Flux.1 and eDiff-I [5] utilize information encoded by both T5 and CLIP to guide image generation. Recently, works like LuminaT2X [19], Kolors [73], and LI-DiT [45] directly use pre-trained decoder-only LLM as the prompt encoder, which enhances the prompt-following ability in image generation. However, these methods can only encode the text prompt, lacking the ability to understand information from images, videos, audio, etc.
@@
MLLMs as Text Encoders for T2I. VLMs [35, 77, 3, 76, 40, 10, 43, 81] have made significant progress in visual language understanding tasks. Furthermore, models such as AnyGPT [91], NextGPT [79], X-LLM [7], and MiniCPM-o [24] have further developed the capabilities in audio understanding. In articles that fuse multimodal models with image generation models, MUMU [6] and ML-MGIE [18] inject multimodal information into the image generation model using VLM but also change the weights of the generation model during training, lacking plug-and-play capabilities. KOSMOS-G [53] and Easyref [95] achieve alignment while preserving the weights of the original SD [16] and SDXL [56] models. KOSMOS-G trains a VLM and AlignerNet from scratch, whereas Easyref trains the features and projector of the final layer output of the VLM. These methods consume significant computational resources and are limited to a single task, failing to fully leverage the zero-shot capabilities of VLM. Furthermore, the aforementioned methods focus on alignment with U-Net-based models like SDXL. Currently, there is a lack of an efficient alignment method for DiT models, such as Flux.1 and SD3. Work in the related field, such as HunyuanVideo [30], requires an additional CLIP text encoder for alignment. Qwen2VL-Flux [44] not only retains the structure of T5, but also alters the weights of Flux.1 itself, significantly reducing the model’s plug-and-play and zero-shot capabilities.
@@
Distillation of DiT. Knowledge distillation (KD) has been widely applied in diffusion models [93, 28, 94]. SSKD [75] direct distillation of information from the crucial attention mechanism from teacher to student can significantly narrow the performance gap between both. EFFICIENT-VDIT [14] adopts a multi-step consistency distillation to accelerate DiT sampling. PEA-Diffusion aligns multilingual CLIP with SDXL using the L2-norm distance, enabling the T2I model to have multilingual capabilities. However, this approach is also designed for the U-Net backbone and consumes considerable computational resources.
Unified Image Generation Models. Numerous works [22, 82, 83, 90, 38] have successively collected billion-level training data to design novel architectures for multi-stage training from scratch. Additionally, there are models [50, 70, 72, 15, 78] based on incremental training of the basic T2I models. All of these models require substantial training resources and data and do not support input modalities beyond textual and visual.
@@ 3.1 Preliminary
where $\mu$ represents Layer Normalization (LN). Then, perform normalization using adaptive parameters $\beta_{2},\gamma_{2}$,
@@ 3.2 Model Overview
The framework consists of two parts, as shown in Fig. 2. The initial part involves X2I training exclusively on a text corpus, with AlignNet as the trainable parameters. Distribution alignment between student and teacher is attained via attention distillation. The subsequent part incorporates the LightControl module, forming X2I Enhanced Training, thereby improving the accuracy of reference images in image instruction editing tasks.
Figure 2: Overview diagram of the X2I. In the deep red box on the left represents the X2I training, where only the text data needs to be input for training, with trainable parameters limited to AlignNet and the distillation location situated within the attention output of MM-DiT. The light green box at the bottom right represents X2I enhanced training after the X2I training, AlignNet is fixed, and the trainable parameters limited to LightControl.
@@ 3.3 AlignNet with MLLM’s Hidden State
To integrate hidden states from all layers of MLLMs, we design a simple CNN to capture spatial and channel-wise relationships between layers. Convolutional operations allow the extraction of more complex patterns and improve the rationality of weight allocation, enabling spatial adaptability and establishing inter-layer dependencies. By adjusting the kernel size, we can also create cross-layer receptive field correlations, achieving better multi-scale fusion.
In MLLM, define a feature extractor that projects text $y$ into a middle feature $H\in\mathbb{H}^{b\times m\times s\times z}$, where $m$ represents all layers of MLLM’s hidden states, each layer has a hidden size of $z$ and a sequence length of $s$. Next, we define a CNN mapping network $\Psi$ to enable self-learning of weights across all layers’ hidden states, resulting in the aggregation of hidden states from $m$ layers into $1$ layers of weighted hidden states as the following formula,
@@ 3.4 Attention Distillation
KD can be divided into two main categories: logits distillation and feature distillation. Logits distillation, also known as response-based KD. In MM-DiT, intermediate layer features encompass low-level information such as pixel values, shape, and gradient information, as well as high-level information such as color, theme, and lighting details. In this paper, the student model is conditioned to inject multimodal feature information. Knowledge transfer occurs layer by layer through the concatenation of the information injected at each level. Therefore, for deeply encoded information that logits cannot express, we opt for intermediate feature distillation between layers to achieve alignment.
Further, each MM-DiT consists of AdaLN as Eq. 1, self-attention layer as Eq. 2, LN as Eq. 3, and FF layers as Eq. 4. In addition to regressing the $\gamma$ and $\beta$ after the LN, AdaLN also regresses a $\alpha$ before the end of each residual module. Since biases here directly affect the distribution of features, we choose to perform feature distillation in the position of the attention output $x_{A},c_{A}$ in order to minimize unnecessary regression and maintain consistency in the distribution affected by normalization parameters. This provides an important conclusion in exploring direct feature distillation in the MM-DiT feature layers as well as the the single-DiT block of Flux.1.
@@
$\mathcal{L}(\theta)=\text{KL}[\o_{l}||\upsilon_{l}]=\left[-\mathbb{E}\log\frac{\upsilon_{l}(A_{t}|\epsilon,c)}{\o_{l}(A_{s}|\epsilon,y)}\right],$ (7)
where ø$\o,\upsilon$ specifically represent the attention output mappings in each layer of the student and teacher models in MM-DiT, respectively. $A_{t}$ denotes the attention output $x_{A},c_{A}$ mentioned above; similarly, $A_{s}$ corresponds to the output of the attention output in the student model. $\epsilon$ represents noise input because we perform distillation during the inference process, which is the reverse denoising process.
@@ 4.2 Tasks and Benchmarks
T2I and Multilingual T2I. Evaluation Datasets. We randomly select five sub-evaluation datasets from T2I-CompBench++[25], each containing 1k images. These include: complex compositions bench, non-spatial bench, attributes binding bench related to color and texture categories, numeracy bench. GenAI-Bench[33], EvalMuse[21], and Multilingual-General[46]. Metrics includes four main categories: image-text matching score indicators like ClipScore (CS) with ViT-bigG[11] and FGA-BLIP2 (FB)[21] Score. Human-consistency feedback score indicators include PickScore (PS)[29], HPSv2[80], and ImageReward (IR)[85]. Fine-grained score indicators like BLIP-VQA (BV)[25] and UniDet[25]. Model scoring indicators reflecting complex textual semantic alignment such as GPT-4o and VQAScore (VS)[39], with GPT-4o instructions specified by the T2I-CompBench++ for complex compositions evaluation. Furthermore, the IR score ranges from 98.2% probability within [-2, 2], and the FB score ranges from [1, 5], we use eight normalized metrics to determine the relative gap compared to the teacher model, defining a Performance Ratio (PR) metric as a performance percentage indicator. Baseline model for universal T2I capability comparison is the Flux.1 teacher model, used to assess reduction in student model’s basic T2I capabilities. Qwen2VL-Flux has similar functions, but MLLM does not support text input and relies solely on the source model’s T5. Thus, it is excluded from comparison. For the comparison of multilingual T2I capabilities, we evaluate both PEA-Diffusion and Sana[84].
@@
ControlNet. Evaluation Datasets is sourced from MultiGen-20M[58] and includes four tasks: A model based on canny edge detection, evaluated using the F1 score. Models based on hed, evaluated using Structure Similarity Index Measure (SSIM). A model based on depth information, evaluated using RMSE.

### v3

@@ Abstract
Text-to-image (T2I) models are well known for their ability to produce highly realistic images, while multimodal large language models (MLLMs) are renowned for their proficiency in understanding and integrating multiple modalities. However, currently there is no straightforward and efficient framework to transfer the multimodal comprehension abilities of MLLMs to T2I models to enable them to understand multimodal inputs. In this paper, we propose the X2I framework, which endows Diffusion Transformer (DiT) models with the capability to comprehend various modalities, including multilingual text, screenshot documents, images, videos, and audio. X2I is trained on a 100K English corpus in 160 GPU hours. Building on the DiT teacher model, we adopt an innovative distillation method to extract the inference capabilities of the teacher model and design a lightweight AlignNet structure to serve as an intermediate bridge. Compared to the teacher model, X2I shows a decrease in performance degradation of less than 1% while gaining various multimodal understanding abilities, including multilingual to image, image to image, image-text to image, video to image, audio to image, and utilizing creative fusion to enhance imagery. Furthermore, it is applicable for LoRA training in the context of image-text to image generation, filling a void in the industry in this area. We further design a simple LightControl to enhance the fidelity of instructional image editing. Finally, extensive experiments demonstrate the effectiveness, efficiency, multifunctionality, and transferability of our X2I. The open-source code and checkpoints for X2I can be found at the following link: https://github.com/OPPO-Mente-Lab/X2I.
@@ 1 Introduction
The text encoder in T2I is critical for generating semantically relevant images, as seen from early models such as CLIP[57] to later models such as T5[58] and advanced LLMs[72, 13]. The main goal is to improve the semantic understanding of text without considering the impact of other modalities on visual output. In contrast, VLMs, including early approaches like Flamingo[2] and BLIP2[35], along with newer models like Llama-Vision[12], QwenVL[3], and InternVL[10], have effectively integrated pre-trained visual encoders with LLMs for enhanced text and visual comprehension. Is there a more streamlined way to transfer VLM or MLLM capabilities to DiT to enable diverse modal inputs? The industry is primarily tackling this challenge through two strategies. One method, similar to PEA-Diffusion[45], utilizes feature distillation by aligning multilingual LLM encoders to the original dimensions using MLP. However, it requires substantial resources of 1600 GPU hours and large-scale image-text datasets. This distillation relies on the U-Net[63] block’s outputs and restricts support for multimodal input. Another method is the VLMs-based replacement of T2I[43], but it is costly, requiring extensive training or fine-tuning of T2I or MLLMs, necessitating substantial data quality and quantity[68, 22]. Alternatively, a solution akin to GlueGen[55] could be employed to align the features of a single-modal or multimodal encoder with the latent space of existing T2I models. However, aligning solely through the encoder still results in suboptimal overall performance.
@@
In this work, we introduce an efficient framework, X2I, along with a simple LightControl module to facilitate the enhancement of image transition from weak to strong fidelity. Utilizing a DiT-based image generation model as a teacher, the student model mirrors the entire architecture of the teacher model but substitutes its text encoder with MLLMs. Both models are capable of processing the same linguistic input. To optimize our training data without gathering images, we opt to distill the teacher model’s inference abilities. In T2I models, this necessitates performing distillation during the reverse denoising phase. Furthermore, we efficiently transfer the visual generation capabilities of the teacher model to the student model by ingeniously designing the AlignNet structure and distilling the attention. As a result, after alignment, the student model acquires the multimodal comprehension abilities of MLLMs during image generation. Moreover, we formulate a template input format intended for the student model to effectively interpret instructions from different modalities. Since MLLMs generally have proficiency in understanding the global semantic features of visual content, we develop LightControl to derive structured data from reference images. This enhances visual fidelity and facilitates precise and controllable image editing.
@@
After training the X2I, we conduct various experiments to compare the capabilities of T2I and instructional-based image editing. Additionally, we conduct subjective experiments for image generation with various multimodal mixed inputs, including multilingual, image (I2I), image-text (IT2I), video (V2I), and audio (A2I). The results demonstrate that X2I acquires robust multimodal understanding capabilities with less than 1% performance degradation in the T2I generation. Furthermore, for more personalized tasks like I2I or IT2I, we conduct experiments involving the training of traditional LoRA[23] to fulfill industry requirements, addressing a training gap in LoRA for these purposes. Moreover, to enhance the adaptability of the framework, only the AlignNet parameters within the model are updated, enabling the aligned student model to accommodate a variety of downstream tasks, such as LoRA, ControlNet[70], IP-Adapter[84], different fine-tuning models, and compression models. Notably, some objective metrics for tasks involving ControlNet and IP-Adapter show improvements over the teacher model. This improvement is due to X2I’s robust support for image conditional information input, which facilitates further extraction of image features. We also perform experimental analyses on training efficiency, attaining 98.2% of the teacher model’s performance with 48K training samples.
In summary, our key contributions are as follows.
@@ 2 Related Work
Diffusion Models and Text Encoder. Diffusion models have shown immense potential in the field of T2I. Earlier works[65, 62, 17, 82] use U-Net as the backbone of the diffusion model, which, to some extent, impacts the scalability of the models. Inspired by DiT[52], many works[8, 32, 17, 36] replace the U-Net backbone with transformers, greatly improving the quality of the generated images. In order to improve the understanding of the diffusion model’s prompts, Stable Diffusion (SD) used CLIP to encode text information, while Flux.1 and eDiff-I[5] utilize information encoded by both T5 and CLIP to guide image generation. Recently, works like LuminaT2X[19], Kolors[71], and LI-DiT[44] directly use pre-trained decoder-only LLMs as the prompt encoder, which enhances the prompt-following ability in image generation. However, these methods can only encode the text prompt, lacking the ability to understand information from images, videos, audio, etc.
@@
MLLMs as Text Encoders for T2I. VLMs[35, 75, 3, 74, 39, 10, 42, 78] have made significant progress in visual language understanding tasks. Furthermore, models such as AnyGPT[87], NextGPT[76], X-LLM[7], and MiniCPM-o[24] have further developed the capabilities in audio understanding. In articles that fuse multimodal models with image generation models, MUMU[6] and ML-MGIE[18] inject multimodal information into the image generation model using VLM but also change the weights of the generation model during training, lacking plug-and-play capabilities. KOSMOS-G[51] and Easyref[91] achieve alignment while preserving the weights of the original SD[16] and SDXL[54] models. KOSMOS-G trains a VLM and AlignerNet from scratch, whereas Easyref trains the features and projector of the final layer output of the VLM. These methods consume significant computational resources and are limited to a single task, failing to fully leverage the zero-shot capabilities of VLM. Furthermore, the aforementioned methods focus on alignment with U-Net-based models like SDXL. Currently, there is a lack of an efficient alignment method for DiT models, such as Flux.1 and SD3. Work in the related field, such as HunyuanVideo[30], requires an additional CLIP text encoder for alignment. Qwen2VL-Flux[43] not only retains the structure of T5 but also alters the weights of Flux.1 itself, significantly reducing the model’s plug-and-play and zero-shot capabilities.
@@
Distillation of DiT. Knowledge distillation (KD) has been widely applied in diffusion models[89, 28, 90]. SSKD[73] involves the direct distillation of critical attention mechanism information from teacher to student, which can considerably reduce the performance gap between both. EFFICIENT-VDIT[14] utilizes a multi-step consistency distillation technique to accelerate DiT sampling. PEA-Diffusion integrates multilingual CLIP with SDXL by employing the L2-norm distance, thus allowing the T2I model to support multilingual capabilities. However, this method is specifically tailored for the U-Net backbone and demands significant computational resources.
@@ 3.1 Preliminary
where $\mu$ represents Layer Normalization (LN). Then, normalization is performed utilizing adaptive parameters $\beta_{2},\gamma_{2}$,
@@ 3.2 Model Overview
The framework is composed of two parts, as shown in Fig. 2. The initial part involves X2I training on a textual dataset, with AlignNet serving as trainable parameters. Distribution alignment between the student and teacher models is achieved through attention distillation. The subsequent part incorporates the LightControl module, forming X2I Enhanced Training, which enhances the precision of reference images in tasks involving image instruction editing.
Figure 2: Overview diagram of the X2I. In the deep red box on the left is the X2I training, where only the text data needs to be input for training, with trainable parameters limited to AlignNet and the distillation location situated within the attention output of $M$-DiT. The light green box at the bottom right represents X2I enhanced training. After the X2I training, AlignNet is fixed, and the trainable parameters are limited to LightControl.
@@ 3.3 AlignNet with MLLM’s Hidden State
To integrate hidden states from various layers of MLLMs, we design a simple CNN to capture spatial and channel-wise interactions across layers. Convolutional operations allow the extraction of intricate patterns and enhance the rationale of weight distribution, permitting spatial flexibility and forming inter-layer linkages. By adjusting the kernel size, we can also generate cross-layer receptive field connections, thus attaining superior multi-scale integration.
In MLLM, define a feature extractor that projects text $y$ into a middle feature $H\in\mathbb{H}^{b\times m\times s\times z}$, where $m$ represents all layers of MLLM’s hidden states, each layer has a hidden size of $z$ and a sequence length of $s$. Next, we define a CNN mapping network $\Psi$ to enable self-learning of weights across all layers’ hidden states, resulting in the aggregation of hidden states from $m$ layers into $1$ layer of weighted hidden states as the following formula,
@@ 3.4 Attention Distillation
KD can be divided into two main categories: logits distillation and feature distillation. In MM-DiT, intermediate layer features encompass low-level information such as pixel values, shape, and gradient information, as well as high-level information such as color, theme, and lighting details. In this paper, the student model is conditioned to inject multimodal feature information. Knowledge transfer occurs layer by layer through the concatenation of the information injected at each level. Therefore, for deeply encoded information that logits cannot express, we opt for intermediate feature distillation between layers to achieve alignment.
Further, each MM-DiT consists of AdaLN as Eq. 1, self-attention layer as Eq. 2, LN as Eq. 3, and FF layers as Eq. 4. In addition to regressing the $\gamma$ and $\beta$ after the LN, AdaLN also regresses an $\alpha$ before the end of each residual module. Since biases here directly affect the distribution of features, we choose to perform feature distillation in the position of the attention output $x_{A},c_{A}$ in order to minimize unnecessary regression and maintain consistency in the distribution affected by normalization parameters. This provides an important conclusion in exploring direct feature distillation in the MM-DiT feature layers as well as the the single-DiT block of Flux.1.
@@
$\mathcal{L}(\theta)=\text{KL}[\eta_{l}||\upsilon_{l}]=\left[-\mathbb{E}\log\frac{\upsilon_{l}(A_{t}|\epsilon,c)}{\eta_{l}(A_{s}|\epsilon,y)}\right],$ (7)
where $\eta,\upsilon$ specifically represent the attention output mappings in each layer of the student and teacher models in MM-DiT, respectively. $A_{t}$ denotes the attention output $x_{A},c_{A}$ mentioned above; similarly, $A_{s}$ corresponds to the output of the attention output in the student model. $\epsilon$ represents noise input because we perform distillation during the inference process, which is the reverse denoising process.
@@ 4.2 Tasks and Benchmarks
T2I and Multilingual T2I. Evaluation Datasets. We randomly select five sub-evaluation datasets from T2I-CompBench++[25], each containing 1k images. These include: complex compositions bench, non-spatial bench, attributes binding bench related to color and texture categories, numeracy bench. GenAI-Bench[33], EvalMuse[21], and Multilingual-General[45]. Metrics includes four main categories: image-text matching score indicators like ClipScore (CS) with ViT-bigG[11] and FGA-BLIP2 (FB)[21] Score. Human-consistency feedback score indicators include PickScore (PS)[29], HPSv2[77], and ImageReward (IR)[81]. Fine-grained score indicators have BLIP-VQA (BV)[25] and UniDet[25]. Model scoring indicators reflecting complex textual semantic alignment such as GPT-4o and VQAScore (VS)[38], with GPT-4o instructions specified by the T2I-CompBench++ for complex compositions evaluation. Furthermore, the IR score ranges from 98.2% probability within [-2, 2], and the FB score ranges from [1, 5]. We use eight normalized metrics to determine the relative gap compared to the teacher model, defining a Performance Ratio (PR) metric as a performance percentage indicator. Baseline model for universal T2I capability comparison is the Flux.1 teacher model, used to assess reduction in student model’s basic T2I capabilities. Qwen2VL-Flux has similar functions, but MLLM does not support text input and relies solely on the source model’s T5. Thus, it is excluded from comparison. For the comparison of multilingual T2I capabilities, we evaluate both PEA-Diffusion and Sana[80].
@@
ControlNet. Evaluation Datasets is sourced from MultiGen-20M[56] and includes four tasks: A model based on canny edge detection, evaluated using the F1 score. Models based on hed, evaluated using Structure Similarity Index Measure (SSIM). A model based on depth information is evaluated using RMSE.